OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (texaschic101):

-3 | 2x + 6 | = -12 ---- divide by -3 | 2x + 6 | = -12/-3 | 2x + 6 | = 4 (2x + 6) = 4 -(2x + 6) = 4 2x = 4 - 6 -2x - 6 = 4 2x = -2 -2x = 4 + 6 x = -1 -2x = 10 x = -5 solutions are x = -1 and x = -5